Sam's age is 4 years more than twice the age of Dave The sum of their ages is 46. What are their ages?

Found 3 solutions by FrankM, Quadratic1600, ikleyn:
Answer by FrankM(1040)

S=2D+4
S+D=46
S=46-D
46-D=2D+4
42=3D
D=14 S=32

Answer by Quadratic1600(28)

Hi
S + D = 46
S = 2D + 4
We now have two simultaneous equations. I am going to say
S = - D + 46
S = 2D + 4
Ok so
46 - D = 2D + 4
46 -4 = 2D + D
42 = 3D
42/3 = D
14 = D
Now we have to work out S
S = 2D + 4

So
S = 2 x 14 + 4
S = 32
Dave is 14
Sam is 32
Hope this helps

Answer by ikleyn(52797)

.
It can be easy solved MENTALLY without using equations.

    Take for a minute these 4 extra years aside.


    Then Sam's "corrected" age will be twice the Dave's age, and the sum of ages will be 46 - 4 = 42.


    Then it is clear that Dave's age is one third of 42, i.e. 14 years, while Sam's "corrected" age is 28 years.


    Now the last step is to return to Sam his 4 years back to get the 


    Answer.  Sam is 32 years old;  Dave is 14 years old.

Completed.